OpenStudy (anonymous):

solve the equation for y 2x+6y=12

OpenStudy (baseballer2014):

2x+6y=12 -2x -2x 6y=12-2x /6 y=2-2/6x y=2-1/3x

OpenStudy (anonymous):

2x+6y=12 subtract 2x 6y=12-2x divide 6 y=12-2x/6

OpenStudy (baseballer2014):

the 2 is a 12 lol sorry

OpenStudy (anonymous):

y=12-2x/6 is the answer